Dustin just retired, and has $530,000 to invest. A very safe Certificate of Deposit (CD) account pays 1.5%, while a riskier bond fund pays 5% in interest. Dustin figures he needs $24,000 a year in interest to live on. How much should he invest in each account to make enough interest while minimizing his risk?

Answer by Boreal(15235)

x with CD and 530000-x for the bond fund
interest earned is .015x+0.05(530000-x)=24000
0.015x+26500-0.05x=24000
-0.035x+26500=24000
-0.035x=-2500
x=$71428.57
x=74128.57 for CD @1.5%=$1071.43 ANSWER
530000-x=$458571.43 for bond fund @5%=$22928.57 ANSWER
That sum is $24,000